### v1

# ReasoningRank: Teaching Student Models to Rank through Reasoning-Based Knowledge Distillation

Abstract
Reranking documents based on their relevance to a given query is a critical task in information retrieval. Traditional reranking methods often focus on improving the initial rankings but tend to lack transparency, failing to explain why one document is ranked higher than another. In this paper, we introduce ReasoningRank, a novel reranking approach that enhances clarity by generating two types of reasoning: explicit reasoning, which explains how a document addresses the query, and comparison reasoning, which justifies the relevance of one document over another. We leverage large language models (LLMs) as teacher models to generate these explanations and distill this knowledge into smaller, more resource-efficient student models. While the student models may not outperform LLMs in speed, they significantly reduce the computational burden by requiring fewer resources, making them more suitable for large-scale or resource-constrained settings. These student models are trained to both generate meaningful reasoning and rerank documents, achieving competitive performance across multiple datasets, including MSMARCO and BRIGHT. Experiments demonstrate that ReasoningRank improves reranking accuracy and provides valuable insights into the decision-making process, offering a structured and interpretable solution for reranking tasks.  [^1] Our code and data are available here https://anonymous.4open.science/r/Distillation-Is-All-You-Need-C98D/README.md
ReasoningRank: Teaching Student Models to Rank through Reasoning-Based Knowledge Distillation
Yuelyu Ji1, Zhuochun Li1, Rui Meng2, Daqing He1 1School of Computing and Information, University of Pittsburgh, 2Salesforce Research {yuj49,zhl163, dah44}@pitt.edu, ruimeng@salesforce.com
1 Introduction
The ability to effectively rank and rerank documents is critical in information retrieval (IR), where the quality of search results significantly impacts user satisfaction Pradeep et al. (2023b); Sun et al. (2023). Traditional reranking approaches often improve initial rankings provided by retrieval systems but pay limited attention to the reasoning behind ranking decisions. Recent studies have demonstrated that incorporating reasoning into the reranking process can enhance performance and interpretability Pradeep et al. (2023b); Liu et al. (2024); Yu et al. (2024c); Sun et al. (2023). By offering explanations that clarify why certain documents are prioritized over others, these methods help users understand and trust the search results.
However, most existing reranking models do not distinguish between two key types of reasoning: explicit reasoning, which explains how a document matches a query, and comparative reasoning, which justifies why one document ranks higher than another. This limitation can reduce ranking effectiveness, particularly in contexts where a subtlety understanding of document relevance is crucial.
Figure 1: Illustration of the two types of reasoning: explicit reasoning provides direct answers to the query, while comparison reasoning evaluates the relative relevance between documents. The LLM generates these explanations to enhance the interpretability of the ranking process.
In this paper, we propose ReasoningRank, a reranking framework that leverages both explicit and comparative reasoning to provide more detailed and transparent document rankings. Our approach is based on knowledge distillation, where a large, powerful teacher model (an LLM) generates high-quality explanations and rankings. These explanations are then distilled into smaller, more computationally efficient student model that perform the reranking tasks. Our contributions are threefold:
  • We introduce a reranking approach that separates explicit and comparative reasoning, enhancing clarity of document rankings.
  • We demonstrate that student models can learn from LLM-generated data, achieving high reranking performance while remaining computationally efficient.
  • We provide an open-source implementation to facilitate further research and applications in information retrieval.
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 presents the proposed ReasoningRank framework. Section 4 not only presents the quantitative performance metrics but also provides a thorough discussion on the implications of the results, analyzing how the reasoning strategies impact the ranking performance and offering insights into the model’s interpretability. Finally, Section 6 concludes the paper and discusses future work.
2 Related Work
The evolution of reranking techniques has significantly advanced with the rise of neural networks, particularly transformers Vaswani (2017); Yu et al. (2024b, a); Zheng et al. (2024); Zhang et al. (2024); Shen et al. (2024); Fan and Tao (2024); Fan et al. (2024). Early models like monoBERT Nogueira et al. (2019) utilized BERT Devlin (2018) in a cross-encoder setup, effectively capturing contextual dependencies between queries and documents. Further models such as monoT5 Nogueira et al. (2020) and InPars Bonifacio et al. (2022) continued to push reranking performance boundaries, especially on datasets like MS MARCO Bajaj et al. (2016). However, these models rely heavily on large-scale fine-tuning and task-specific data, limiting their scalability and adaptability to new domains or tasks.
Figure 2: Overview of explicit and comparison reasoning prompts for document ranking. Explicit reasoning explains how a document matches a query, while comparison reasoning evaluates the relative relevance between documents.
Figure 3: Overview of the ReasoningRank framework. The teacher model generates explicit and comparative reasoning, which is used to train student models capable of reproducing the reranked order and generating explanations.
Early ranking systems primarily focused on pointwise methods, where each document’s relevance was judged in isolation Liu et al. (2009). However, pairwise approaches like duoBERT Pradeep et al. (2021) were introduced to compare pairs of documents, determining which document was more relevant. While this improved the relative ranking of documents, it still missed the larger picture of the entire ranked list. To address this, listwise models such as RankGPT Sun et al. (2023) and its distilled versions like RankVicuna Pradeep et al. (2023a) considered the entire ranked list during training, allowing for more holistic improvements in ranking accuracy. However, even these methods have limitations when it comes to explaining why a particular document ranks higher than another.
The Learning to Rank framework introduced several foundational approaches in information retrieval, including pointwise, pairwise, and listwise methods, which form the basis of modern neural ranking models Liu et al. (2009). Pointwise approaches focus on predicting the relevance score of individual documents in isolation, while pairwise methods aim to indicate the relative ranking between document pairs. On the other hand, listwise approaches focus on directly optimizing the entire ranking list, offering better overall accuracy by considering the global structure of the ranking. Despite these advances, most early methods lacked clarity that modern applications demand.
Recent work has also highlighted the importance of natural language explanations in improving clarity and trust in reranking decisions. Models such as Explain-then-Rank Zhuang et al. (2023) have focused on generating explanations that describe how a document matches a query. However, these approaches often only address explicit reasoning (i.e., document-query matching) and do not account for comparative rationale (i.e., why one document ranks higher). Our work builds on these efforts by incorporating both types of reasoning, enhancing the overall clarity and coherence of the explanations. This dual focus provides more precise insights into the ranking process and improves the model’s ability to deliver accurate and logically consistent rankings Doshi-Velez and Kim (2017); Lipton (2018); Mo et al. (2024); Yu et al. (2024b, a).
Knowledge distillation has become an essential technique in training smaller, more efficient student models from larger teacher models Hinton et al. (2015) to further reduce the computational demands of large reranking models. Knowledge distillation has been particularly effective in reranking tasks, as demonstrated by models like RankVicuna Pradeep et al. (2023a), which distilled the reranking capabilities of RankGPT into a smaller, more efficient model without significant loss of performance. Our approach builds on this by applying distillation to both the reasoning and ranking components, ensuring that our student models not only maintain high performance but also offer explainability in their reranking decisions Qin et al. (2023); Dai et al. (2022); Xu et al. (2024b, a).
By integrating these advancements, our ReasoningRank framework improves upon existing methods by offering a dual reasoning strategy—explicit and comparative reasoning—within a computationally efficient model. Our approach not only enhances ranking accuracy but also provides clear and easy-to-understand explanations for the ranking decisions, addressing fundamental limitations in performance seen in prior work.
3 Method
The overview of the architecture of the ReasoningRank model is shown in Figure 3. We propose ReasoningRank, a framework that integrates explicit and comparative reasoning within a unified reranking approach.
3.1 Teacher Model with Dual Reasoning
Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the teacher model processes these inputs to generate two types of reasoning and rerank order. The teacher model can generate high-quality reasoning that combines both explicit and comparative aspects, and the examples can be seen in Figure 2: Explicit Reasoning: The model explains how each document directly addresses the query, focusing on the relevance and specificity of the content. Comparative Reasoning: The model assesses the relative relevance between documents, explaining why one document should be ranked higher based on content.
3.2 Student Model with Knowledge Distillation
Deploying large language models (LLMs) for reranking is often impractical due to their high computational cost. To address this, we designed the student model to replicate the teacher’s performance while being more efficient for real-world deployment. The innovation lies in using knowledge distillation: the student learns from the teacher’s detailed reasoning and ranking outputs, retaining accuracy and interpretability with significantly lower computational demands. This approach ensures the student can perform complex reranking tasks efficiently, making it ideal for large-scale use.
Architecture and Training:
Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the model processes these inputs to generate both reasoning outputs and ranking scores.
The model generates two outputs: Reasoning Outputs: The model generates textual reasons $\mathbf{r}_{i}$ explaining the relevance of each document $d_{i}$ to the query $q$. Ranking Scores: The model simultaneously generates ranking scores $\mathbf{s}_{i}$ for each document, indicating its relevance to the query.
3.2.1 Training Objective and Loss Functions
To optimize the performance of the student model, we define three distinct loss functions, each corresponding to different aspects of the model. First, using pairwise loss to optimize the relative ranking between pairs of documents, ensuring that the more relevant document is ranked higher. The pairwise loss is defined as:
$\mathcal{L}_{\text{pairwise}}=\sum_{(i,j)\in P}\max(0,1-(\mathbf{s}_{i}-\mathbf{s}_{j}))$
where $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ represent the predicted relevance scores of documents $d_{i}$ and $d_{j}$, and $P$ is the set of document pairs with known relevance relations.
Next, the listwise loss optimizes the overall ranking of the document list. It uses Kullback-Leibler (KL) divergence to measure the difference between the predicted ranking distribution and the target ranking distribution:
$\mathcal{L}_{\text{listwise}}=D_{\text{KL}}(Q(\mathbf{z})\parallel P(\mathbf{s}))$
Where $P(\mathbf{s})=\text{softmax}(\mathbf{s}_{i})$ is the predicted ranking distribution based on the model’s output, and $Q(\mathbf{z})=\text{softmax}(\mathbf{z}_{i})$ is the target ranking distribution based on true labels.
Finally, the generation loss optimizes the quality of the textual reasoning generated by the model. Based on the cross-entropy loss between the generated reasoning and the ground-truth reasoning:
$\mathcal{L}_{\text{generation}}=-\sum_{i=1}^{n}\mathbf{y}_{i}^{\text{gen}}\log(\text{softmax}(\mathbf{r}_{i}))$
where $\mathbf{r}_{i}$ represents the generated reasoning logits for document $d_{i}$, and $\mathbf{y}_{i}^{\text{gen}}$ is the ground truth reasoning label.
3.3 Final Training Objective
To achieve balanced training, we introduce learnable weights $\alpha$, $\beta$, and $\gamma$ to adjust the contributions of pairwise, listwise, and generation losses, respectively. The final loss function is:
$\mathcal{L}=\alpha\times\mathcal{L}_{\text{pairwise}}+\beta\times\mathcal{L}_{\text{listwise}}+\gamma\times\mathcal{L}_{\text{generation}}$
Where $\alpha+\beta+\gamma=1$, ensuring a balanced combination of the three loss functions during training. With this design, the student model efficiently performs ranking tasks and provides transparent reasoning for its decisions, enhancing both transparency and performance.
4 Experiments and Results
In this section, we evaluate the performance of our proposed ReasoningRank framework across various information retrieval tasks. We introduce the datasets and evaluation metrics used in our experiments (§4.1). We then detail the experimental setup and baseline methods for comparison (§4.3). Finally, we present a comprehensive analysis of the experimental results, including comparisons with existing methods in (§4.4). We performed paired t-tests across all queries in each dataset to assess the statistical significance of the improvements. The annotations in Tables 1 and 2 indicate the significance levels.
4.1 Datasets and Evaluation Metrics
In addition to MSMARCO Bajaj et al. (2016) and BEIRThakur et al. (2021), we utilize the BRIGHT Su et al. (2024) to further evaluate our model’s ability to handle reasoning-intensive retrieval tasks.
Evaluation Metrics
We use the following evaluation metrics to measure the performance of our models:
Normalized Discounted Cumulative Gain (NDCG@k): Measures the ranking quality of the retrieved documents, with higher emphasis on top-ranked documents. We report NDCG at rank positions 5 and 10.
BLEU, ROUGE-L, and Coherence Score: Used for evaluating the quality of generated reasons in reasoning tasks, assessing both lexical overlap and logical consistency.
4.2 Teacher Models
We utilize large language models (LLMs) as teacher models to generate reasoning annotations: GPT-4: A state-of-the-art LLM known for its strong reasoning and language understanding capabilities. ClaudeThe : An advanced LLM developed by Anthropic, offering competitive performance in reasoning tasks, we used the version claude-3-5-sonnet-20240620. GeminiTeam et al. (2023): A commercial LLM used for comparison purposes.We used the version gemini-1.5-flash.
4.2.1 Overall Performance on Benchmark Datasets
Table 1 summarizes the NDCG@5 performance of different models GPT-4 consistently achieves the best results, with 77.69% on DL19 and 85.27% on TREC-COVID when using both comparison and explicit reasoning prompts, showcasing its robust handling of complex queries.
Adding different reasoning prompts to the LLMs enhances their performance across various datasets. For instance, GPT-4 with comparison and explicit reasoning outperforms its baseline (without reasoning) on DL19 by 2.10%. Similar trends are observed with Claude and Gemini, indicating that incorporating reasoning prompts in different LLMs is beneficial.
To confirm the statistical significance of the improvements, we conducted paired t-tests comparing models with reasoning prompts to their respective baselines. The results, indicated by and in Table 1, show that the enhancements are significant, with p-values less than 0.05.
Given GPT-4’s overall solid performance and the boost from reasoning prompts, we select it as the teacher model to distill its enhanced reasoning abilities into our student models.
Models DL19 DL20 Covid Touche News NFCorpus Robust04 DBPedia
BM25 50.58 47.96 59.47 44.22 39.52 30.75 40.73 31.83
Rank GPT-4 Models
Rank GPT-4 (without reasoning) 75.59 70.56 85.51 38.47 57.55 36.62 57.55 47.12
+ Explicit reasoning 75.32 69.93 83.27 38.94 55.87 36.08 52.89 46.98
+ Comparison reasoning 76.25 71.07 84.62 38.23 56.49 36.52 56.53 48.15
+ Comparison & Explicit 77.69 73.24 85.27 38.29 58.43 36.33 58.63 49.52
Claude Models
Claude (without reasoning) 71.96 68.81 82.77 36.12 54.35 36.36 50.78 42.51
+ Explicit reasoning 70.59 68.23 82.22 35.92 55.43 35.59 52.38 42.51
+ Comparison reasoning 71.34 69.75 83.31 36.43 55.11 36.71 51.53 43.23
+ Comparison & Explicit 72.13 69.95 84.03 37.33 55.52 36.36 52.67 44.91
Gemini Models
Gemini (without reasoning) 71.10 68.45 82.95 35.62 53.65 35.49 50.12 42.92
+ Explicit reasoning 69.08 67.02 81.55 34.83 52.87 36.92 51.31 43.92
+ Comparison reasoning 70.17 68.24 82.41 35.91 53.35 35.42 51.12 42.85
+ Comparison & Explicit 71.35 68.57 83.47 36.97 53.80 36.12 51.78 43.67
Table 1: NDCG@5 performance (in percentage) for different models with various reasoning prompts across multiple datasets. Bold numbers indicate the best performance within each group of models. indicates statistically significant improvement over the base model without reasoning prompts (p $<$ 0.05). indicates statistically significant improvement over both the base model and the single reasoning prompt variants (p $<$ 0.05).
Models DL19 DL20 Covid Touche News NFCorpus Robust04 DBPedia
Student Models
Basic w/o reasoning 73.23 70.80 79.94 35.10 51.90 34.60 52.82 46.31
With explicit reasoning 74.54 70.11 79.37 34.25 52.81 35.45 50.75 46.95
With comparison reasoning 74.12 72.25 80.10 34.00 53.10 36.63 52.11 47.21
Comparison & Explicit 75.39 72.36 84.59 36.22 53.81 36.37 53.54 47.85
Baseline Models
DeBERTa Sun et al. (2023) 68.53 64.18 73.42 32.13 50.24 33.67 49.23 45.39
MonoT5 Nogueira et al. (2020) 74.54 70.44 79.95 34.07 46.03 35.21
RankVicuna Pradeep et al. (2023a) 71.12 68.66 67.14 48.73 38.53 55.67 35.29
RankZephyr Pradeep et al. (2023b) 72.23 70.53 85.12 36.46 53.33 38.92 60.72 35.52
APEER Jin et al. (2024) 74.57 72.25 83.86 35.25 52.11 33.41 56.03 46.13
Table 2: NDCG@5 performance (in percentage) for student models and baseline comparisons across multiple datasets. indicates statistically significant improvement over the baseline DeBERTa model (p $<$ 0.05). Bold indicates the best performance among all models for each dataset. Entries marked with "–" indicate that results are not available for that dataset.
4.3 Student Models
Experimental Setup
We implement our models using LoRA (Low-Rank Adaptation) on 32GB V100 GPUs to fine-tune the student models on large datasets efficiently. Based on the LLaMA 3.1 8B architecture, the student models fine-tuned with different reasoning prompts derived from the teacher models.
4.3.1 Performance Comparison with Baseline
Our student models, trained with explicit and comparison reasoning, achieve competitive performance compared to the previous state-of-the-art student models. As shown in Table 2,the improvements over the baseline models are statistically significant, with p-values less than 0.05 (indicated by ).
BM25: A traditional term-based retrieval model using TF-IDF, known for its simplicity and efficiency in keyword-based retrieval.
RankGPT DeBERTa Sun et al. (2023): A transformer model using disentangled attention to better capture long-range dependencies, improving reranking by understanding query-document context.
MonoT5 Nogueira et al. (2020): A T5-based model treating reranking as a sequence generation task, leveraging its encoder-decoder architecture for better relevance scoring.
RankVicuna, RankZephyr and APEER Pradeep et al. (2023a, b); Jin et al. (2024): RankVicuna is a distilled version of RankGPT, designed for zero-shot listwise reranking with fewer parameters, balancing efficiency and performance.RankZephyr a state-of-the-art zero-shot listwise reranking model that adapts RankGPT prompts for robust performance with minimal task-specific training. APEER share the same structure with RankZephyr but have an automated prompt generation for large language models.
4.4 ReasoningRank Student Model vs. Baseliness
Our ReasoningRank student model demonstrates strong performance across various datasets, often outperforming the baseline models. For instance, on the DL19 dataset, our student model achieves an NDCG@5 of 75.39%, surpassing DeBERTa (68.53%), RankVicuna (71.12%), and RankZephyr (72.23%). Similarly, on the DL20 dataset, our model attains 72.36%, outperforming DeBERTa (64.18%) and RankVicuna (68.66%).
On some datasets, our model achieves the best performance among all models. For example, on the News dataset, our student model reaches an NDCG@5 of 53.81%, slightly outperforming RankZephyr (53.33%) and significantly better than DeBERTa (50.24%).
However, there are cases where other models perform better. On the Touche dataset, RankVicuna achieves an NDCG@5 of 48.73%, which is higher than our student model’s 36.22%. This may be due to RankVicuna’s design for zero-shot listwise reranking, which could be particularly effective for argument retrieval tasks represented in Touche. Similarly, on the Robust04 dataset, RankZephyr attains a higher NDCG@5 of 60.72% compared to our model’s 53.50%.
These variations suggest that while our model performs well, certain models may have advantages in specific domains or datasets due to their architectures or training strategies. For example, RankVicuna’s strong performance on Touche might be attributed to its effectiveness in handling argumentation and persuasive texts, which are prevalent in that dataset. RankZephyr’s higher score on Robust04 could be due to its ability to leverage zero-shot capabilities effectively in that particular dataset.
4.4.1 Performance on Reasoning-Intensive Tasks (BRIGHT Dataset)
The BRIGHT dataset is specifically designed to evaluate models on reasoning-intensive retrieval tasks, simulating real-world scenarios where complex reasoning is required to determine document relevance. Table 3 shows the NDCG@5 results on the BRIGHT dataset. In the Earth Science domain, the student model achieves an NDCG@5 of 27.0%, closely approaching the teacher model’s 31.0%. Similarly, in Psychology, the student model attains 26.0%, compared to the teacher’s 28.9% and BM25’s 11.6%.
Domain BM25 Teacher Student GPT-4 COT
Bio. 17.2 24.3 21.0 28.0
Earth. 24.2 31.0 27.0 37.1
Econ. 13.2 19.0 16.5 22.1
Psy. 11.6 28.9 26.0 34.2
Rob. 12.0 16.0 14.5 19.0
Stack 16.0 21.0 18.5 24.2
Sus. 12.4 18.0 16.2 21.9
Leet. 22.1 27.6 25.0 31.4
Pony 8.5 5.9 5.2 6.7
AoPS 6.0 7.6 6.8 9.2
TheoQ. 9.4 18.5 15.3 23.0
TheoT. 3.5 17.0 14.2 21.7
Average 13.1 19.6 17.5 23.2
Table 3: NDCG@5 results on the BRIGHT dataset.
4.4.2 Quality of Generated Reasons
We assess the quality of the generated reasons using BLEU and ROUGE-L as presented in Table 4 compared with the teacher model’s reason. The baseline model here represents the model trained without the rank loss function and has the same input as ours. On the DL19 dataset, the student model achieves a BLEU score of 21.4, compared to the baseline’s 18.0, and a ROUGE-L score of 36.8 versus 32.5. Similar improvements are observed on DL20, where the student model attains BLEU and ROUGE-L scores of 24.0 and 38.2, respectively, surpassing the baseline’s 19.5 and 34.0.
Dataset Model BLEU ROUGE-L
DL19 Baseline 18.0 32.5
Ours 21.4 36.8
DL20 Baseline 19.5 34.0
Ours 24.0 38.2
Covid Baseline 16.0 30.0
Ours 18.9 34.5
Touche Baseline 17.5 33.0
Ours 20.5 37.1
News Baseline 20.0 35.0
Ours 22.7 39.3
Table 4: Evaluation of reason generation using BLEU, and ROUGE-L scores for baseline and student models.
4.4.3 Ablation Study: Impact of Reasoning Strategies
Results
In Table 5, by removing comparison reasoning results in a performance drop on DL19 from 75.39 to 74.12 (NDCG@5). Similarly, removing explicit reasoning causes a performance decrease on DL20 from 72.36 to 72.25 (NDCG@5).
Analysis
Both reasoning strategies contribute to performance improvements. The ablation study shows that omitting either reasoning component results in a noticeable decline in performance, highlighting their complementary roles in improving ranking effectiveness.
Model DL19 NDCG@5/@10 DL20 NDCG@5/@10
Comparison & explicit 75.39 / 73.85 72.36 / 71.22
w/ explicit reasoning 74.54 / 73.18 70.11 / 70.97
w/ comparison reasoning 74.12 / 72.94 72.25 / 70.43
Basic w/o reasoning 73.23 / 72.12 70.82 / 69.83
Table 5: Ablation study: Impact of explicit and comparison reasoning on NDCG@5/@10 for DL19 and DL20.
4.4.4 Effects of Training Data Size Across Student Models
Figure 4 i shows the performance of small models trained on 100, 1,000, and 2,000 examples. Contrary to expectations, increasing data size from 1,000 to 2,000 does not consistently improve performance and sometimes slightly reduces it. We observed similar trends across various small models, including Mistral and LLaMA variants, highlighting that adding more data does not always guarantee better performance.
Figure 4: Student model performance for different training data sizes on DL19 and DL20.
5 Discussion
Our experiments demonstrate that the ReasoningRank framework effectively enhances document reranking by incorporating explicit and comparative reasoning. The student models achieve competitive results with significantly reduced computational resources compared to large teacher models. The statistically significant improvements confirm the effectiveness of reasoning strategies in enhancing model performance.
An important aspect of ReasoningRank is its ability to not only rank documents but also generate coherent and relevant reasoning for its rankings. This dual capability suggests that the model could be applied to other tasks that require both ranking and reasoning generation.
While our current work focuses on document reranking in information retrieval, the approach of integrating reasoning generation with ranking decisions could be extended to other domains. For example, in recommendation systems, the model could rank products or services while providing explanations for its recommendations, thereby improving user trust. Similarly, in question answering or summarization tasks, generating reasoning alongside answers or summaries could enhance understandabilityJiang et al. (2024).
However, we acknowledge that our paper does not evaluate the model’s generative abilities on tasks beyond document reranking. This is a limitation in assessing the full scope of the model’s applicability. Future work should explore how the model performs in different contexts to verify its generalizabilityMcElfresh et al. (2022); Rahmani et al. (2022).
6 Conclusion
In this paper, we introduced ReasoningRank, a novel approach to document reranking that incorporates explicit and comparative reasoning. Our method leverages the power of large language models to generate interpretable reasons, which are then distilled into smaller, more efficient student models. Our experiments demonstrate the effectiveness of this approach, achieving high performance across multiple datasets while maintaining interpretability.
Limitations
While our approach shows strong performance, it has several limitations. First, the reliance on large language models for generating reasoning limits the applicability of our method in resource-constrained environments. Second, the quality of the generated reasons heavily depends on the quality of the prompts, which requires careful crafting. Future research should address these limitations by exploring more efficient models and automated prompt-generation techniques.
Ethics Statement
This work enhances the transparency and accessibility of reasoning capabilities in AI by distilling knowledge from LLMs to smaller models. While this improves efficiency, we acknowledge the ethical concerns of bias in LLMs and the need for careful mitigation in real-world applications.
Appendix A Student Training Parameters
Hyperparameter Value
low_cpu_fsdp True Interpretability
run_validation True
batch_size_training 1
context_length 1024
gradient_accumulation_steps 1
gradient_clipping True
gradient_clipping_threshold 1.0
num_epochs 3
max_train_step 0
max_eval_step 0
learning rate 1e-4
weight_decay 0.0
seed 42
use_fp16 True
use_peft True
freeze_layers False
Table 6: Training configuration hyperparameters.
Appendix B Teacher Model Parameters
Hyperparameter Value
Model GPT
Temperature 1.0
Top p 0.9
Table 7: Teacher model hyperparameter settings.
Appendix C NDCG@5 for Different Models
Table 8: NDCG@5 for different models and training data sizes on various datasets.
Model Train Size DL 19 DL 20
Mistral v0.1 100 69.69 76.33
1000 71.03 77.06
2000 70.63 73.99
Mistral v0.3 100 69.26 79.55
1000 71.56 79.14
2000 69.86 76.96
LLaMA 3.1 8B 100 69.63 75.85
1000 72.87 77.03
2000 73.37 77.60
LLaMA 3 8B 100 69.57 77.15
1000 73.76 75.35
2000 71.75 75.55
LLaMA 2 7B 100 71.50 79.89
1000 72.19 72.95
2000 72.18 75.36
Appendix D Model Behavior Analysis
In the reasoning generation tasks, the performance of the model varied across explicit reasoning, comparison reasoning, and their combination. Below, we provide an analysis of unexpected behavior based on 2,000 queries.
Repetition.
The repetition metric measures the occurrence of duplicate passage identifiers. Across the tasks, we observed:
  • Explicit reasoning: 63 duplicate passage identifiers (3%).
  • Comparison reasoning: 36 duplicate passage identifiers (1.5%).
  • Combined reasoning: 47 duplicate passage identifiers (2%).
Missing Documents.
Missing documents were most frequent in explicit reasoning:
  • Explicit reasoning: 642 missing documents (32% of queries).
  • Comparison reasoning: 227 missing documents (11% of queries).
  • Combined reasoning: 328 missing documents (16% of queries).
Appendix E API Cost Analysis
Utilizing large language models (LLMs) like GPT-4 as teacher models involves significant computational and monetary costs. In this section, we analyze these costs based on our experiments.
We calculated the average number of tokens, the number of API requests, and the estimated cost per query for different operations involving the teacher model. The costs are based on the pricing provided by the API service at the time of our experiments 9.
API Instruction Tokens Cost (\$USD)
gpt-4 Explicit 3650 0.134
gpt-4 Comparison 4050 0.158
gpt-4 Comparison & Explicit 4650 0.194
Table 9: Average token usage and cost per query using GPT-4 for different reasoning tasks.
Appendix F Prompts for Reasoning
Table 10: Basic prompt, explicit reasoning prompt, and comparison reasoning prompt.
Basic Prompt:
I will provide you with num passages, each indicated by a number identifier [].
For each passage, briefly generate your reasoning process as follows:
1. Judge whether the passage is not applicable (not very common), where the query does not meet the
premise of the passage.
2. Check if the query contains direct evidence. If so, judge whether the query meets or does not meet
the passage.
3. If there is no direct evidence, try to infer from existing evidence and answer one question:
If the passage is ranked in this order, is it possible that a good passage will miss such information?
If impossible, then you can assume that the passage should not be ranked in that order.
Otherwise, it should be ranked in that order.
Explicit Reasoning Prompt:
Then, read every passage one-by-one and find the sentence where the method is the direct answer
for the query and output the sentence. Then, rank the passages based on their relevance to the query.
Give highest priority to passages that provide a clear and direct definition or explanation of the
keywords related to the query.
Consider both the detailed information and any relevant background context provided in each passage.
Provide clear and concise reasons for the ranking, highlighting the specific parts of the passages that
influenced your decision.
Make sure to ignore any irrelevant information and focus on content directly related to the query.
In addition to direct reasons for ranking each passage, also consider listwise reasons.
Comparison Reasoning Prompt:
A listwise reason involves comparing passages with each other to determine their relative importance.
Specifically, compare passages to identify which ones provide more comprehensive, relevant, and accurate
information in relation to the query.
When providing listwise reasons, mention specific comparative insights, such as why one passage might be
more relevant than another based on the overall context and detail provided.
Return Type:
Search Query: query. Rank the num passages above based on their relevance to the search query and
the extracted keywords. The passages should be listed in descending order using identifiers.
The most relevant passages should be listed first. The output should be in JSON format.
For each passage, provide the extracted keywords and detailed reasons for its ranking.
Ensure to separate direct reasons and listwise reasons clearly.
Mention specific parts of the passage that influenced your decision.
Ensure the reasons are clear, concise, and directly related to the query,
balancing direct definitions or explanations and relevant background information.
Only output the JSON structured format.